OpenStudy (anonymous):

slove for x. x^2+50=10x

OpenStudy (anonymous):

This does not factor. Do know how to complete the square?

OpenStudy (anonymous):

Or do you want to use the quadratic formula? Do you care which way?

OpenStudy (anonymous):

I don't care, I just need to solve for x and I cannot figure it out. Thank you.

OpenStudy (anonymous):

Set the equation equal to 0 first \[x^2-10x+50=0\] To complete the square subtract the constant to the RHS; square half the -10; and add to both sides: \[x^2-10x+25=50+25\] Factor the LHS as a perfect square trinomial: \[(x-5)^2=75\] Square root both sides and add 5 to both sides: \[x = 5\pm 5 \sqrt{3}\] These are your two answers.

OpenStudy (anonymous):

Thank you!